# Latent Beam Diffusion Models for Decoding Image Sequences

### v1

@@ Abstract
While diffusion models excel at generating high-quality images from text prompts, they struggle with visual consistency in image sequences. Existing methods generate each image independently, leading to disjointed narratives — a challenge further exacerbated in non-linear storytelling, where scenes must connect beyond adjacent frames. We introduce a novel beam search strategy for latent space exploration, enabling conditional generation of full image sequences with beam search decoding. Unlike prior approaches that use fixed latent priors, our method dynamically searches for an optimal sequence of latent representations, ensuring coherent visual transitions. To address beam search’s quadratic complexity, we integrate a cross-attention mechanism that efficiently scores search paths and enables pruning, prioritizing alignment with both textual prompts and visual context. Human evaluations confirm that our approach outperforms baseline methods, producing full sequences with superior coherence, visual continuity, and textual alignment. By bridging advances in search optimization and latent space refinement, this work sets a new standard for structured image sequence generation.
@@ 1 Introduction
Figure 1: BeamDiffusion models leverage search in the latent denoising space for the best sequence of denoised images, where $I_{n}$ represents the image at step $n$ and $S_{n}$ represents the text associated with step $n$.
@@
Image diffusion models [42, 12, 43] have made significant advancements in generating high-quality images. These models, especially when combined with text prompts, have shown great potential in producing detailed and contextually accurate visuals [34, 29, 37]. While they excel at generating individual images based on specific text prompts, they face challenges when it comes to creating coherent sequences of images. In most cases, each image is generated independently based on its own prompt, which does not naturally ensure visual continuity across multiple steps. This problem is further compounded when the narrative is non-linear, where a scene may be connected not only to the immediate prior scene but also to earlier scenes in the sequence. This dynamic can result in sequences that lack coherence and visual consistency over time [4].
@@
We propose to utilize beam search as a potential solution to this problem. Beam search is a well-established technique in text generation tasks like machine translation [18, 47, 21, 36], allowing the model to explore multiple possible paths during the generation process, evaluating different trajectories and refining its predictions. However, despite its success in other domains, it has been barely explored in our context, leaving its potential untapped. In full image sequence generation, beam search could help navigate the complex latent space, refining coherence by adjusting text prompts or adding semantic cues.
@@
In this work, we explore beam search for latent prior selection in image sequence generation. While previous research has used latent representations as priors to improve sequence consistency [4, 32], it remains unclear which specific representations yield the best overall sequence. We build on these ideas by introducing beam search to explore the latent space, enabling iterative refinement of image sequences. As illustrated in Figure 1, our approach leverages search in the latent space to identify the best sequence of images, minimizing the risk of falling into suboptimal or incoherent paths, while maintaining visual continuity across images.
Due to the quadratic complexity of beam search algorithm, one needs to prune the search paths to keep a fixed number of candidate paths. To this end, we incorporate a cross-attention mechanism [23] that evaluates and scores candidate image paths based on their alignment with both the textual description of the next step and the visual context of previous steps. This provides a reliable quality measure for each path, allowing us to retain the best ones while pruning the least suitable, ensuring both adherence to instructions and visual continuity throughout the sequence.
@@
We evaluated the effectiveness of our approach on three datasets: in the Recipes domain, and in the out-of-domain DIY Tasks and Visual Storytelling. Human evaluations show that our model outperformed baseline methods for generated sequence quality, showcasing its adaptability to diverse contexts.
2 Related Work
Decoding methods are essential for enhancing the quality of large language models (LLMs). Techniques such as beam search, top-k and nucleus sampling are widely studied for their effectiveness in balancing computational cost with generating coherent, contextually appropriate outputs [40, 13, 9]. Beam search has also proven effective in tasks like machine translation [27], image captioning [45, 19, 1], and visual storytelling [15]. Building on these ideas, our work innovates by applying beam search to explore the latent space in diffusion models for image sequence generation.
Ensuring coherence remains a significant challenge in multimodal sequence generation [10, 24, 6, 20]. Some approaches, like AR-LDM [28], encode the context of caption-image pairs into multimodal representations to maintain consistency across outputs. Similarly, Rahman et al. [31] incorporates the history of U-net [35, 38] latent vectors to preserve narrative coherence. However, these methods often face challenges such as computational inefficiency and the risk of noise propagation over time. Other approaches, such as Gill [22], combine LLMs with image encoder-decoder systems. While effective for certain tasks, these models suffer from alignment issues, with retrieved or generated images failing to match the narrative context, and they too often require complex training. Our method avoids these pitfalls by direct conditional generation against the guiding text while explicitly maintaining coherence via the beam search.
Interpreting diffusion models has become a focal point in recent research. Pont-Tuset et al. [30] anchor each word in textual descriptions to specific regions of images. Tang et al. [44] analyze text-image interactions through cross-attention maps in the denoising module, offering insights into how textual prompts influence the generated images. Dewan et al. [7] enhance this interpretability by leveraging partial information decomposition to analyze how various components of a text prompt impact the generated image. In contrast, our beam-search approach focuses on generating images directly from textual input. Rather than mapping text to pre-existing images, it generates image sequences that are visually coherent and are consistently aligned with the evolving sequence of textual narrative.
Bordalo et al. [4] propose to use latent information from previous steps to guide coherent generation of subsequent images. However, their method relied on a greedy heuristic-based selection. CoSeD [32] improved on this idea and introduced a contrastive learning approach [5] to rank images generated from different latent seeds. Yet, both approaches focused on local refinements rather than optimizing across the entire sequence. Our beam search approach lifts these local selection criteria to perform global optimization.
3 Synthesis of Non-Linear Visual Sequences
In the following section, we define the challenges in image sequence generation, recall the core principles of LDM, Sec. 3.1, and then we detail BeamDiffusion key aspects in terms of diffusion beams (Sec. 3.2.1) and pruning (Sec. 3.2.2) with a contrastive approach (Sec. 3.2.3).
3.1 Problem Formulation
Generating a full sequence of visually coherent images from textual scene descriptions presents a challenge. Given a sequence of scenes $S=\{s_{1},s_{2},\dots,s_{L}\}$, where each scene $s_{j}$ is a textual description, our goal is to generate a corresponding sequence of images $\{I_{1},I_{2},\dots,I_{L}\}$ that maintain consistency across all steps.
Latent Diffusion Models (LDMs) [34] generate high-quality images in a compressed latent space. Instead of directly operating in pixel space, LDMs encode an input image $I$ into a latent representation $z$ using an encoder $\mathcal{E}$ and reconstruct it through a decoder $\mathcal{D}$ [2]. The U-Net backbone [35], denoted as $\epsilon_{\theta}$, incorporates a cross-attention mechanism [17] to condition the diffusion process on an external input $\tau_{\theta}(s)$, where $s$ represents a textual prompt. LDMs are trained by minimizing:
$L_{\text{LDM}}=\mathbb{E}_{\mathcal{E}(I),s,\epsilon\thicksim\mathcal{N}(0,1),t}\left[\|\epsilon-\epsilon_{\theta}(z_{t},t,\tau_{\theta}(s))\|_{2}^{2}\right].$ (1)
While effective, LDMs generate images independently, starting from a randomly initialized latent $z_{T}$. This lack of temporal coherence makes it challenging to ensure consistency across generated sequences, especially in non-linear narratives [41], where maintaining character identity, lighting, and environmental details are crucial.
@@ 3.2 Beam Diffusion Model
To address the limitations of standard LDMs in generating coherent image sequences, we introduce the BeamDiffusion model. The proposed approach aims to enhance both the consistency and quality of image sequences by adopting a strategy inspired by beam search, but with a twist, it operates on latent denoising space. BeamDiffusion incorporates latent representations $\mathcal{Z}_{j}$ from previously generated images. These latent representations guide the generation of new candidates, ensuring that each new image aligns with the context of the evolving sequence. This approach allows visual elements from previous images to be carried over or referenced, maintaining contextual consistency and continuity, while also adapting to the new scene description. In the following sections, we will detail the proposed method.
@@ 3.2.1 Latent Space Exploration with Diffusion Beams
To maintain both diversity and coherence in the generated image sequence, we propose a diffusion-based exploration of the latent space across multiple steps. At each generation step $j$, the set $\mathcal{I}_{j}=\{I_{j}^{(1)},I_{j}^{(2)},\dots,I_{j}^{(B)}\}$ consists of all possible candidate images, with $I_{j}^{(b)}$ representing the image generated at step $j$ for beam $b$.
@@
In the first step of the sequence, $s_{1}$, the model generates multiple candidate images using different random seeds. This approach allows for a broader exploration of the latent space and reduces Bayes risk [3], as described in Appendix B. Hence, in this initial step, $s_{1}$, we obtain the first exploration samples from the denoising latent space, denoted as $\mathcal{Z}_{1}=\{z_{1}^{(1)},\ldots,z_{1}^{(K)}\}$.
Building on prior work [4, 32], incorporating latent denoising from earlier steps ($s_{<j}$) can enhance the denoising process. We define $\mathcal{Z}_{j}$ to be the accumulation of latent representations from all preceding images up to step $j$:
$\mathcal{Z}_{j}=\bigcup_{i\in\{j-n,\ldots,j-1\}}\{z_{i}^{(1)},\dots,z_{i}^{(K)}\},$ (2)
where each latent representation $z_{i}^{(k)}$ corresponds to the $k$-th iteration of the diffusion process for image $I_{i}$. By utilizing latents from $n$ previous stages (steps back), our approach ensures contextual consistency across generations.
Considering one particular search beam $b$ in the latent space, to generate a new candidate image $I_{j}^{(b)}$, we condition the diffusion process on a latent representation $z$ selected from the accumulated latents in $\mathcal{Z}_{j}$, as represented in Figure 2(a). This conditional dependency is captured by the following probability expression:
$\small p\left(I_{j}^{(b)}\mid\theta,z\in\mathcal{Z}_{j},s_{j}\right),$ (3)
where $\theta$ are the model parameters. This formulation ensures that all denoising processes are correlated, as each new image $I_{j}^{(b)}$ is generated based on latents $z$ drawn from $\mathcal{Z}_{j}$. In this way, each iteration in the denoising process acts as a seed for the next, establishing a dependency between successive denoising processes (Figure 2(b)).
@@
Following the idea of beam search, which maintains a set of the most promising candidates at each step and extends them sequentially [46], we construct diffusion beams by iteratively expanding candidate images. Formally, for each step $j$, we generate the set of all candidate images $\mathcal{I}_{j}$ to compute the set of candidate diffusion beams $\hat{B}_{j}$. At the initial step $j=1$, there are no previous beams to combine, so the initial set of beams can be represented as,
$\small\hat{B}_{1}=\bigcup_{k}\{I_{1}^{(k)}\mid I_{1}^{(k)}\in\mathcal{I}_{1}\},$ (4)
where $\mathcal{I}_{1}$ is the set of initial candidate images generated using different random seeds indexed by $k$. For subsequent steps $j>1$, each beam $\hat{B}_{j-1}^{(b)}$ from the previous step is extended with new candidate images $I_{j}^{(b\cdot k)}$ from the set $\mathcal{I}_{j}$, and this extension is represented by the operator $\oplus$. The new set of beams is now represented as,
$\small\hat{B}_{j}=\bigcup_{(b,k)}\left\{\hat{B}_{j-1}^{(b)}\oplus I_{j}^{(b\cdot k)}\mid I_{j}^{(b\cdot k)}\in\mathcal{I}_{j}\right\}$ (5)
where $b$ is the index of a beam from the set $\hat{B}_{j-1}$, and $k$ is the $k$-th latent variable from $\mathcal{Z}_{j}$. For each combination of a previous beam $\hat{B}_{j-1}^{(b)}$ and latent $\mathcal{Z}_{j}^{(k)}$, the model considers a new candidate image $I_{j}^{(b\cdot k)}$ from the set of candidate images $\mathcal{I}_{j}$. The union over $(b,k)$ combines all possible sequences of previous beams with the new candidate images, expanding the search space for the next step of generation.
@@ 3.2.2 Pruning Diffusion Beam Paths
Tracking diffusion beams helps maintain context and coherence across the growing image sequence by exploring multiple potential continuations. However, as the sequence progresses, focusing on the most promising paths within the beam budget (the number of beams that can be maintained simultaneously) becomes increasingly important to ensure the coherence of the generated images and optimize computational resources. To achieve this, our model prunes the diffusion beam paths, narrowing down the search space for the best candidates while promoting diversity and consistency in the maintained beams.
To prune candidate beams, the model scores each candidate $\hat{B}_{j}$ by the likelihood of image $I_{j}$ in the beam given the previous images $I_{<j}$ and the step text. This is represented as $p(I_{j}|s_{j},B_{j-1}^{(b)})$, where $B_{j-1}$ represents the set of the most promising candidate beams from the previous step. Formally, the scoring is defined as:
$\small score(\hat{B}_{j}^{(b)})=\sum_{i=1}^{j}\log p(I_{i}|I_{1},\dots,I_{i-1},s_{1},\ldots,s_{i}),$ (6)
where we sum the individual log-probabilities of each image in the sequence, reflecting the overall coherence and alignment with the context set by the previous images and prompts.
After scoring all candidate images in $\mathcal{I}_{j}$, the search space is refined by performing a beam search operation. This operation selects the most promising $w$ diffusion beams, effectively limiting the beam width to $w$. The refinement process is expressed as,
$\small B_{j}=\mathop{\mathrm{argmax}}_{B_{j}\subseteq\hat{B}_{j},|B_{j}|=w}\left(\sum_{i=1}^{w}score(\hat{B}_{j}^{(i)})\right)$ (7)
where, $B_{j}$ represents the most promising $w$ candidate diffusion beams at the step $j$, ensuring that the updated beam evolves to explore the most promising options, guided by the previous context and the newly generated candidates, while maintaining a controlled beam width.
To maintain a broad search space in the early stages, pruning is delayed until after the second step. During this initial phase, all candidates are retained for exploration. After the second step, pruning is applied, limiting the candidates to the $w$ most promising. This approach balances exploration and computational efficiency, focusing the model’s efforts on the most promising candidates in the later stages of generation.
Figure 3: Visualization of cross-attention in latent diffusion, showing how BeamDiffusion focuses on different regions of the image based on the prompt at each stage. The heatmap highlights the areas with the strongest attention.
3.2.3 Contrastive Score
While the basic pruning mechanism outlined above provides a foundation for selecting candidate images based on raw probabilities, the scoring function is now implemented as contrastive classifier. This classifier improves the image selection by learning from past sequences and aligning the candidates with the evolving context [32]. As illustrated in Figure 3, the heat maps show how BeamDiffusion attends to different regions of the input text at various stages, highlighting the evolving alignment between textual prompts and image features.
To compute these probabilities, we refine the approach using a softmax-based method to determine the likelihood of each candidate image $I_{j}$ relative to all other possible candidates in the sequence. In this refined process, the probability $p(I_{j}|s_{j},B_{j-1}^{(b)})$ is replaced by the output of the classifier $\varphi(I_{j},s_{j},B_{j-1}^{(b)})$, which approximates the likelihood of each candidate fitting the sequence. The more detailed version of the scoring mechanism can be expressed as,
$\scriptsize\begin{split}score(\hat{B}_{j}^{(b)})&=\sum_{i=1}^{j}\log\left(\frac{\exp\left(\varphi(I_{i},s_{i},B_{i-1}^{(b)})\right)}{\sum_{I_{i}^{\prime}\in\mathcal{I}_{i}}\exp\left(\varphi(I_{i}^{\prime},s_{i},B_{i-1}^{(b)})\right)}\right),\end{split}$ (8)
where the denominator in the softmax function normalizes the scores by summing over the candidate images $I_{i}^{\prime}$ for beam $b$ within the set $\mathcal{I}_{i}$. This ensures that the score of selecting a particular candidate image $I_{j}$ is computed relative to all other candidate images in the sequence, capturing how well it fits the overall sequence in terms of visual content and textual alignment.

### v2

@@ Abstract
While diffusion models excel at generating high-quality images from text prompts, they struggle with visual consistency in image sequences. Existing methods generate each image independently, leading to disjointed narratives — a challenge further exacerbated in non-linear storytelling, where scenes must connect beyond adjacent frames. We introduce a novel beam search strategy for latent space exploration, enabling conditional generation of full image sequences with beam search decoding. Unlike prior approaches that use fixed latent priors, our method dynamically searches for an optimal sequence of latent representations, ensuring coherent visual transitions. As the latent denoising space is explored, the beam search graph is pruned with a cross-attention mechanism that efficiently scores search paths, prioritizing alignment with both textual prompts and visual context. Human and automatic evaluations confirm that BeamDiffusion outperforms other baseline methods, producing full sequences with superior coherence, visual continuity, and textual alignment. All data and code implementations will be released after publication.
@@ 1 Introduction
Image diffusion models [43, 9, 44] have made significant advancements in generating high-quality images. These models, especially when combined with text prompts, have shown great potential in producing detailed and contextually accurate visuals [35, 29, 38]. Despite the impressive results at generating individual images based on specific text prompts, diffusion models face challenges when it comes to creating coherent sequences of images. In most cases, each image is generated independently based on its own prompt, which does not naturally ensure visual continuity across a sequence of correlated prompts. This problem is further compounded when the narrative is non-linear, where a scene may be connected not only to the immediate prior scene but also to earlier scenes in the sequence. This dynamic can result in sequences that lack coherence and visual consistency over time [2].
@@
We propose to utilize beam search as a potential solution to this problem. Beam search is a well-established technique in text generation tasks like machine translation [14, 51, 17, 37], allowing the model to explore multiple possible paths during the generation process, evaluating different trajectories and refining its predictions. However, despite its success in other domains, it has been barely explored in this context, leaving its potential untapped. In full image sequence generation, beam search could help navigate the complex latent space, refining coherence by adjusting to multiple text prompts or adding semantic cues.
@@
In this work, we explore beam search for latent prior selection in image sequence generation. While previous research has used latent representations as priors to improve sequence consistency [2, 33], it remains unclear which specific representations yield the best overall sequence. We build on these ideas by introducing beam search to explore the latent space, enabling iterative refinement of image sequences. As illustrated in Figure 1(a), our approach leverages search in the latent space to identify the best sequence of images, minimizing the risk of falling into suboptimal or incoherent paths, while maintaining visual continuity across images. As the beam diffusion algorithm traverses the latent space, the number of hypothesis grow and the least probably search paths are pruned to keep a fixed number of candidate paths. To this end, we incorporate a cross-attention mechanism [19] that evaluates and scores candidate image paths based on their alignment with both the textual description of the next step and the visual context of previous steps. This provides a reliable quality measure for each path, allowing us to retain the best ones while pruning the least suitable, ensuring both adherence to instructions and visual continuity throughout the sequence.
@@
We evaluated the effectiveness of BeamDiffusion in three domains: the Recipes domain and in the out-of-domain DIY and Visual Storytelling. Results obtained with human annotations, automatic evaluations and LLM-as-a-judge, all showed that the proposed approach outperforms baseline methods for generated sequence quality, showcasing its adaptability to diverse contexts.
@@
2 Related Work
Decoding methods are crucial for enhancing the quality of large language models (LLMs). Techniques such as beam search, top-k and nucleus sampling are widely studied for their effectiveness in balancing computational cost with generating coherent, contextually appropriate outputs [40, 10, 7]. Beam search has also proven effective in tasks like machine translation [25], image captioning [48, 15], and visual storytelling [12].
Ensuring coherence remains a major challenge in multimodal sequence generation [8, 22, 4, 16]. Some approaches, like AR-LDM [28], encode the context of caption-image pairs into multimodal representations to maintain consistency across outputs. Likewise, Rahman et al. [32] incorporates the history of U-net [36, 39] latent vectors to preserve narrative coherence. Other approaches, such as Gill [18], combine LLMs with image encoder-decoder systems. StoryDiffusion [52], changes the U-Net’s cross-attention mechanism enabling training-free consistency by sampling tokens from previous images to maintain coherence in newly generated frames. StoryGen [21] denoises the previous image and the resulting features serve as conditioning for the subsequent generation of images. GenHowTo [45] uses a control net during inference to project the previous image directly into the latent space to guide the diffusion process for the next image step. StackedDiffusion [26] processes stacked image sequences as a single latent representation allowing the attention mechanism to consider all steps together, enhancing coherence across steps. In [2], authors proposed to use latent information from previous steps to guide coherent generation of subsequent images relying on a greedy heuristic-based selection. CoSeD [33] improved on this idea and introduced a contrastive learning approach [3] to rank images generated from different latent seeds. However, both approaches focused on local refinements rather than optimizing throughout the sequence. Our beam search approach extends these local selection criteria to perform global optimization.
While effective for certain tasks, these models suffer from alignment issues, with retrieved or generated images failing to match the narrative context, and they too often require complex training. Our method addresses these previous constraints by generating images directly from the guiding text, maintaining consistency via a beam search for optimal seeds, similar to methods such as [41] that view image generation as a search process.
3 Generating Non-Linear Visual Sequences
Generating a full sequence of visually coherent images from textual scene descriptions is a challenging research problem. Given a sequence of steps $S=\{s_{1},s_{2},\dots,s_{L}\}$, where each sequence step $s_{j}$ is a textual description, our goal is to generate a corresponding sequence of images $\{I_{1},I_{2},\dots,I_{L}\}$ that maintain consistency across all steps of the sequence. While Latent Diffusion Models (LDMs) [35] can generate high-quality images in a compressed latent space, this latent space is designed to generate images independently, starting from a randomly initialized latent $z_{T}$.
LDMs encode an input image $I$ into thelatent space representation $z$ using an encoder $\mathcal{E}$ and reconstruct it through a decoder $\mathcal{D}$ [35]. The reverse diffusion process is modeled by a U-Net backbone [36], $\epsilon_{\theta}$, which incorporates a cross-attention mechanism [47] to condition the denoising iterations on an external input, e.g. the output embeddings of a text encoder $\tau_{\theta}(s)$. LDMs are trained to iteratively estimate a denoised variant of $z$, with the loss
$L_{\text{LDM}}=\mathbb{E}_{\mathcal{E}(I),s,\epsilon\sim\mathcal{N}(0,1),t}\left[\|\epsilon-\epsilon_{\theta}(z_{t},t,\tau_{\theta}(s))\|_{2}^{2}\right]$ (1)
This loss evidences the lack of awareness of other denoising processes, making it challenging to ensure consistency acrosscorrelated prompts or previous denoising processes. This is particularly challenging in non-linear narratives [42], where maintaining character identity, lighting, and environmental details are crucial. To address these challenges, prior work [32, 21, 45, 2, 33] reuse latent denoising iterations from earlier steps in the sequence, paving the way to a key question: how to determine the optimal set of denoising latents that maximizes coherence across different generations? The answer lies in a combinatorial search space, that BeamDiffusion can traverse with a search strategy to determine an optimal solution, as we will discuss next.
@@ 3.1 Beam Diffusion Model
To address the limitations of standard LDMs in generating coherent image sequences, we introduce the BeamDiffusion model. The proposed approach aims to enhance both the consistency and quality of image sequences by adopting a strategy inspired by beam search, but with a twist, it operates on latent denoising space. BeamDiffusion incorporates latent representations from previously generated images. These latent representations guide the generation of new candidates, ensuring that each new image aligns with the context of the evolving sequence. This approach allows visual elements from previous images to be carried over or referenced, maintaining contextual consistency and continuity, while also adapting to the next scene description.
In the remainder of this article, the term step will always refer to an element in the input sequence; and the term iteration will refer to one denoising iteration of the diffusion model.
@@ 3.1.1 Latent Space Exploration with Diffusion Beams
To generate coherent and diverse image sequences, we explore the latent space of diffusion models using a beam search-inspired strategy. For each sequence step $j$ we generate a set of candidate images $\mathcal{I}_{j}=\{I_{j}^{(1)},I_{j}^{(2)},\dots\}$, where each $I_{j}^{(k)}$ is conditioned on a different beam path.
Algorithm 1 BeamDiffusion Algorithm
1:  Input parameters: $\{s_{1},\ldots,s_{N}\}$: sequence of steps; $r$: number of random seeds; $m$: steps back; $w$: beam width.
2:  $\hat{B}_{1}\leftarrow\{\text{SD}(s_{1},rnd)\}^{r}$triangleright Beam initialization with $r$ different images.
3:  for $j=2$ to $L$ do
4:     for all beam $b\in\hat{B}_{j-1}$ do
5:        $\mathcal{Z}_{j}\leftarrow$$\{\hat{B}_{j-n},\ldots,\hat{B}_{j-1}\}$triangleright Read denoising latents from previous steps.
6:        for all latent $z\in\mathcal{Z}_{j}$ do
7:           $I_{j}\leftarrow\text{SD}(s_{j},z)$triangleright Generate candidate a image.
8:           $\hat{B}_{j-1}\leftarrow\hat{B}_{j-1}\bigcup{I_{j}}$triangleright Extend beam with candidate image.
9:        end for
10:     end for
11:     $\text{score}(I_{j})\leftarrow\varphi(I_{j}\in\hat{B}_{j})$triangleright Score the new image of the beam.
12:     $B_{j}\leftarrow\arg\max_{|B_{j}|=w}\text{score}(\hat{B}_{j})$triangleright Prune beams and keep top $w$ beams.
13:  end for
14:  return  best beam $B^{*}\leftarrow\arg\max_{B_{L}}\text{score}(B_{L})$
@@
In our approach, the denoising process for each sequence step is conditioned by representations from prior steps, helping the model maintain coherence across the sequence. In the first step of the sequence, $s_{1}$, we generate multiple candidate images by running the diffusion process with $r$ random noise seeds (Alg. 1, line 2). This strategy enables broader exploration of the latent space and reduces Bayes risk [1], as detailed in Appendix D.
For every generation process corresponding to a sequence step $s_{i}$, we store the first $N$ latent space vectors $z_{i}=\{z_{i,t}\}_{t\in\subseteq\{1,\ldots,N\}}$ produced during the denoising iterations. On each step prompt $s_{j>1}$ we need to generate diverse samples that (a) explore different denoising trajectories in the latent space, and (b) are strongly correlated with previous samples. Leveraging the stored latent space vectors, we sample the latent subspace of $n$ previous denoising processes. Hence, for steps $j>1$, to enable contextual consistency, we first accumulate representations from the previous $m$ steps:
$\mathcal{Z}_{j}=\bigcup_{i\in\{j-m,\ldots,j-1\}}\{z_{i}\}.$ (2)
To generate new candidate images $\mathcal{I}_{j}=\{I_{j}^{(1)},\ldots,I_{j}^{(k)},\ldots,I_{j}^{(N\cdot m)}\}$ at step $j$, we condition the diffusion model on the latent trajectories in $\mathcal{Z}_{j}$. Specifically, for each latent trajectory $z\in\mathcal{Z}_{j}$, the model performs a new diffusion process to produce a corresponding image candidate $I_{j}^{(k)}$, Alg. 1, line 5 and 7. This step establishes a dependency between generations, where latents from earlier steps guide the synthesis of visually coherent transitions over time, as illustrated in Figure 1.
@@
Inspired by beam search [50], which maintains and extends the most promising candidates at each sequence step, we introduce diffusion beams to structure sequential image generation. At each step $j$, the set of candidate images $\mathcal{I}_{j}$ is used to compute the set of candidate diffusion beams $\hat{B}_{j}$: for each step $j$ we consider all the beams existing in step $j-1$ and extend each one of such $b$ beams $\hat{B}_{j-1}^{(b)}$ by pairing it with a new candidate image $I_{j}^{(k)}\in\mathcal{I}_{j}$. The resulting beams are defined as
$\hat{B}_{j}=\bigcup_{(b,k)}\left\{\hat{B}_{j-1}^{(b)}\oplus I_{j}^{(b\cdot k)}\mid I_{j}^{(b\cdot k)}\in\mathcal{I}_{j}\right\},$ (3)
where $\oplus$ denotes the concatenation of a prior beam with a new candidate image, and $(b\cdot k)$ indexes the pairing of beam $b$ and candidate image $I_{j}^{(k)}$. This process builds a tree of candidate sequences, Alg. 1, line 8, enabling the model to explore diverse yet coherent visual narratives over time.
@@ 3.1.2 Pruning Diffusion Beam Paths
Tracking multiple diffusion beams enables BeamDiffusion to maintain visual coherence while exploring diverse continuations across a sequence. However, as the sequence length increases, the number of candidate beams grows rapidly, making it computationally infeasible to retain them all. To address this, we apply a pruning strategy based on a learned contrastive scoring function, which allows us to retain only the most promising beams while preserving diversity.
At each sequence step $j$, we construct a set of candidate beams $\hat{B}_{j}$. Each candidate beam $\hat{B}_{j}^{(b)}$ is formed by extending a beam from the previous step $\hat{B}_{j-1}^{(b)}$ with a newly generated image $I_{j}^{(k)}$. A complete candidate beam at step $j$ is a sequence of images $\hat{B}_{j}^{(b)}=[I_{1}^{(b)},I_{2}^{(b)},\ldots,I_{j}^{(b)}]$, one per step. To evaluate the overall quality of a beam, we use a contrastive classifier [33], denoted by $\varphi$. Specifically, for each image $I_{i}^{(b)}$ in the beam, the classifier $\varphi$ computes a compatibility score based on the image’s prompt $s_{i}$ and the previous images in the beam $B_{i-1}^{(b)}$, which includes earlier images and prompts in the same beam. The total beam score is then the sum of these per-image scores over all sequence steps up to $j$:
$score(\hat{B}_{j}^{(b)})=\sum_{i=1}^{j}\log\left(\frac{\exp\left(\varphi(I_{i}^{(b)},s_{i},B_{i-1}^{(b)})\right)}{\sum_{I_{i}^{\prime}\in\mathcal{I}_{i}}\exp\left(\varphi(I_{i}^{\prime},s_{i},B_{i-1}^{(b)})\right)}\right),$ (4)
where $\mathcal{I}_{i}$ is the set of all candidate images at sequence step $i$. The contrastive formulation ensures that each image is evaluated not in isolation, but relative to other candidates, capturing how well it fits the evolving sequence context.
Once scored, we prune the beam candidates by selecting the top $w$ scoring beams:
$B_{j}=\mathop{\mathrm{argmax}}_{B_{j}\subseteq\hat{B}_{j},|B_{j}|=w}\left(\sum_{i=1}^{w}score(\hat{B}_{j}^{(i)})\right),$ (5)
where $B_{j}$ denotes the final set of retained beams for sequence step $j$.
To balance exploration and efficiency, pruning is disabled for the initial steps, allowing all candidates to be retained initially. After a few step, we apply pruning to keep only the top $w$ beams. This strategy ensures wide exploration early on and focused, high-quality continuation in later sequence steps.